OpenStudy (anonymous):

The bases are trapezoids with bases of lengths 7cm and 9cm perpendicular to one side of length 6cm, and the height of the prism is 12cm.

OpenStudy (anonymous):

area of base\[=\frac{ 7+9 }{ 2 }*6=48 cm^2\] \[volume =base~area*height=48*12~ cm^3=?\]

OpenStudy (anonymous):

thanks!!